The decomposition of nitramide in aqueous solution at 25 °C

NH2NO2(aq)------->N2O(g) + H2O(l)

is first order in NH2NO2.

During one experiment it was found that when the initial concentration of NH2NO2 was 0.458 M, the concentration of NH2NO2 dropped to 0.117 M after 21860 s had passed.

Based on this experiment, the rate constant for the reaction is _________s-1.
Answer #1

we have:

[NH2NO2]o = 0.458 M

[NH2NO2] = 0.117 M

t = 21860.0 s

use integrated rate law for 1st order reaction

ln[NH2NO2] = ln[NH2NO2]o - k*t

ln(0.117) = ln(0.458) - k*2.186*10^4

-2.146 = -0.7809 - k*2.186*10^4

k*2.186*10^4 = 1.365

k = 6.243*10^-5 s-1

Answer: 6.24*10^-5 s-1